Solve the system by substitution.
2x + 4y = 10
y = 2x + 15

Answer by sdmmadam@yahoo.com(530)

Solve the system by substitution.
2x + 4y = 10 ----(1)
y = 2x + 15 ----(2)
Putting (2) in (1) (that is substituting for y)
Then (1) becomes
2x + 4(2x+15) = 10
2x +(8x +60) = 10
(2x+8x) =10-60 [By additive associativity and adding (-60) to both the sides]
10x = -50
x = (-50)/10 = -5
Putting x=-5 in (2),
y = 2x +15 becomes
y = 2X(-5) +15 = -10 + 15 = +5
Answer: x = -5 and y = 5
Verification: putting x = -5 and y = 5 in (1)
LHS = 2x + 4y = 2X(-5)+ 4X(5) = -10+20 = 10 =RHS
Therefore our answer is correct.